40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 25 20:44:11 2012, Jenne, Update, SUS, Tip tilts in progress 
    Reply  Fri Oct 26 10:44:25 2012, Jenne, Update, SUS, Tip tilts in progress 
       Reply  Fri Oct 26 13:08:14 2012, Jenne, Update, SUS, Tip tilts in progress 
Message ID: 7630     Entry time: Fri Oct 26 10:44:25 2012     In reply to: 7625     Reply to this: 7631
Author: Jenne 
Type: Update 
Category: SUS 
Subject: Tip tilts in progress 

Quote:

Jamie and I spent some time with tip tilt SN001 this afternoon.  This was installed as SR3, so I was going to put a new LaserOptik mirror in there.  I accidentally snapped one of the wires (I forgot how strong the magnets are - one zipped from the mirror holder and captured the wire).  Jamie and I put the new LaserOptik mirror in, with the wedge correct, but we need to re-resuspend it with the 0.0036" wire tomorrow.  We'll also keep working on re-pitch aligning the other optics.

PR2 needs to be put back as a G&H, and we need to put a LaserOptik mirror into PR3.

 We resuspended SN001 this morning with 0.0036" wire.  We did as Koji suggested, and flipped the wire clamp so the suspension point is a little higher, so we'll see if that helps.  We put LaserOptik mirror SN1 into this TT001.

We put the G&H mirror back into TT004, which is PR2.  We also put a LaserOptik mirror (SN5) into TT005, which is SR3.

Jamie is working on re-pitch aligning TT004 and TT005 (we already did 001), then we can re-install them in the vacuum system later this afternoon.

ELOG V3.1.3-